Price increases from $10 to $12 and the price elasticity of demand is -0.5. The quantity demanded is 450 units. What was it before?


1
Expert's answer
2021-12-17T08:40:30-0500

"E=\frac{\frac{Q2-Q1}{Q1}}{\frac{P2-P1}{P1}}"

"-0.5=\frac{\frac{450-Q1}{Q1}}{\frac{12-10}{10}}"

Q1=500